# M2PT: Multimodal Prompt Tuning for Zero-shot Instruction Learning

### v2

@@ Abstract
Taowen Wang1thanks: Equal contribution., Yiyang Liu [^1] footnotemark: 1, James Chenhao Liang1, Junhan Zhao2, Yiming Cui3, Yuning Mao4, Shaoliang Nie4, Jiahao Liu 5, Fuli Feng6, Zenglin Xu7,8, Cheng Han9, Lifu Huang10, Qifan Wang4, Dongfang Liu1 1Rochester Institute of Technology, 2Harvard Medical School, 3ByteDance, 4Meta AI, 5Meituan, 6University of Science and Technology of China, 7Shanghai Academy of AI for Science, 8Fudan University, 9University of Missouri - Kansas City, 10University of California - Davis {tw9146, dongfang.liu}@rit.edu
@@ 4.1 Experiment Setup
For training, we conduct instruction tuning on Vision-Flan Xu et al. (2024b), which is a human-annotated multimodal instruction tuning dataset with $191$ diverse tasks. To reduce computational costs, we follow common practice Shen et al. (2024) and employ a scaled-down version containing up to $1,000$ instances per task, resulting in a total of $191,105$ instances. For zero-shot evaluation, we examine our approach on the comprehensive multimodal evaluation benchmark MME Yin et al. (2023), measuring both perception and cognition abilities across $14$ subtasks. We further evaluate the model’s capabilities using $7$ multimodal datasets. Specifically, for Optical Character Recognition, we utilize the Text-VQA Singh et al. (2019), and for reasoning, we employ the Visual Spatial Reasoning (VSRLiu et al. (2023a). Following Zhai et al. (2023); Shen et al. (2024), the perception capability is tested on CIFAR-10/100 Krizhevsky et al. (2009) and MNIST Deng (2012). SNLI-VE Xie et al. (2019) evaluates Visual Entailment capabilities, while the POPE Li et al. (2023) dataset examines the tendency towards object hallucination. More details are provided in the Appendix S1.
@@ 5 Analysis and Discussion
Following common practice Sun et al. (2024b), we extract and discuss the activation maps from the attention block of MLLMs, and investigate the influence of visual and textual prompts in Fig. 3. We randomly select two samples from the MME dataset and visualize their corresponding activation maps of the attention block in the last layers of both visual encoder (Fig. 3(a)) and LLM (Fig. 3(b)).
To analyze the impact of textual and visual prompts to the frozen components, we categorize them, according to LLaVa’s model structure, into textual prompts, system text tokens, visual prompts, image tokens and instructions. Several findings can be observed.
@@
We further conduct failure case studies for M2PT on CIFAR-10 dataset. The results in Fig. 9 indicate that the predominant misclassification within the CIFAR-10 (i.e., 44% or 443 examples of all such misclassification) is the mis-identification of “ships” as “boats”. This high error rate could potentially be attributed to the limited resolution of the images within the CIFAR-10, coupled with the inherent similarity in the semantic characteristics between “ship” and “boat”. Additional case studies and discussions are provided in Appendix S4.
@@ 7 Acknowledgements
This research was supported by the National Science Foundation under Grant No. 2242243.

### v4

@@ Abstract
Taowen Wang1thanks: Equal contribution., Yiyang Liu [^1] footnotemark: 1, James Chenhao Liang1,11, Junhan Zhao2, Yiming Cui3, Yuning Mao4, Shaoliang Nie4, Jiahao Liu 5, Fuli Feng6, Zenglin Xu7,8, Cheng Han9, Lifu Huang10, Qifan Wang4, Dongfang Liu1 1Rochester Institute of Technology, 2Harvard Medical School, 3ByteDance, 4Meta AI, 5Meituan, 6University of Science and Technology of China, 7Shanghai Academy of AI for Science, 8Fudan University, 9University of Missouri - Kansas City, 10University of California - Davis 11U.S. Naval Research Laboratory {tw9146, dongfang.liu}@rit.edu
@@ 4.1 Experiment Setup
For training, we conduct instruction tuning on Vision-Flan Xu et al. (2024b), which is a human-annotated multimodal instruction tuning dataset with $191$ diverse tasks. To reduce computational costs, we follow common practice Shen et al. (2024) and employ a scaled-down version containing up to $1,000$ instances per task, resulting in a total of $191,105$ instances. For zero-shot evaluation, we examine our approach on the comprehensive multimodal evaluation benchmark MME Fu et al. (2023), measuring both perception and cognition abilities across $14$ subtasks. We further evaluate the model’s capabilities using $7$ multimodal datasets. Specifically, for Optical Character Recognition, we utilize the Text-VQA Singh et al. (2019), and for reasoning, we employ the Visual Spatial Reasoning (VSRLiu et al. (2023a). Following Zhai et al. (2023); Shen et al. (2024), the perception capability is tested on CIFAR-10/100 Krizhevsky et al. (2009) and MNIST Deng (2012). SNLI-VE Xie et al. (2019) evaluates Visual Entailment capabilities, while the POPE Li et al. (2023) dataset examines the tendency towards object hallucination. More details are provided in the Appendix S1.
@@ 5 Analysis and Discussion
Following common practice Sun et al. (2024b), we extract and discuss the activation maps from the attention block of MLLMs, and investigate the influence of visual and textual prompts in Fig. 3. We randomly select two samples from the MME dataset and visualize their corresponding activation maps of the attention block in the last layers of both visual encoder (Fig. 3(a)) and LLM (Fig. 3(b)). To analyze the impact of textual and visual prompts to the frozen components, we categorize them, according to LLaVa’s model structure, into textual prompts, system text tokens, visual prompts, image tokens and instructions. Several findings can be observed.
@@
We further conduct failure case studies for M2PT on CIFAR-10 dataset. The results in Fig. 9 indicate that the predominant misclassification within the CIFAR-10 (i.e., 44% or 443 examples of all such misclassification) is the misidentification of “ships” as “boats”. This high error rate could potentially be attributed to the limited resolution of the images within the CIFAR-10, coupled with the inherent similarity in the semantic characteristics between “ship” and “boat”. Additional case studies and discussions are provided in Appendix S4.
@@ 7 Acknowledgements
This research was supported by the National Science Foundation under Grant No. 2242243. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of U.S. Naval Research Laboratory (NRL) or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.